A reaction vessel contains NH3, N2, and H2 at equilibrium at a certain temperature.

The equilibrium

concentrations are [N2] = 0.61 M, [H2] = 1.26 M, and [NH3] = 0.30 M.

Calculate the

equilibrium constant, Kc, if the reaction is represented as

(

1

2

)

N2(g) +

(

3

2

)

H2(g) ⇌ NH3(g)
Answer #1

The reaction that occurs is:

N2 + 3 H2 = 2 NH3

The constant is calculated:

K = [NH3] ^ 2 / [N2] * [H2] ^ 3 = (0.3) ^ 2 / 0.61 * (1.26) ^ 3 = 0.07

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.